ihavewrizz:

Tyler is building a fish tank in the shape of a sphere that has a diameter of 20 inches. Based on Tyler's research, it is recommended to have no more than 1 goldfish for every 693 cubic inches of water. How many goldfish should Tyler buy for his fish tank? Numerical Answers Expected!

ihavewrizz:

@supie @hero

ihavewrizz:

@sxdsouls

ihavewrizz:

@luigi0210

Luigi0210:

Okay so the volume of a sphere is shown as \[\large V= \frac{4}{3} \pi r^3 \] The diameter is given as 20 in, to get the radius, just halve it, so r= 10 in \[\large V = \frac{4}{3} \pi (10)^3 \] Giving you \[\large V= 4188.80 in^3 \] To get the number of goldfish, just divide by the 693 in recommended.

Luigi0210:

So \[\large \frac{ 4188.80}{693} = 6.0 fish\]

ihavewrizz:

ok tysm